OpenStudy (btaylor):

You are travelling in a spaceship with a velocity of +100 m/s. On the bottom of the spaceship is a tennis ball gun, pointing in the direction opposite of the direction you are going. This gun fires a tennis ball with a velocity of -50 m/s. To a static observer, what would the motion of the ball be?

OpenStudy (btaylor):

I think it would appear to travel 'forward' (in the same direction as you), with a velocity of +50 m/s. Is this right? @TuringTest @myininaya

OpenStudy (anonymous):

Nop....@BTaylor......it would appear as the ball comming with a velocity of 150 m/s

OpenStudy (anonymous):

Vs = 100 m/s vb = -50 m/s vs - vb = 150 m/s

OpenStudy (shane_b):

I agree with @BTaylor. Picture the ship moving to the right at 100m/s and the ball being shot to the left at 50m/s.

OpenStudy (anonymous):

@experimentX wat do u think..

OpenStudy (dominusscholae):

@yahoo right concept but its 100 + (-50) = 50 m/s upward. The velocity for a static observer is the vector sum of the velocities, not the subtraction.

OpenStudy (dominusscholae):

The negative is referring to the fact you're shooting it downward with respect to the ground.

OpenStudy (anonymous):

but @dominusscholae u see.... if we take... space ship direction as + ve.... then opposite will be -ve

OpenStudy (anonymous):

since the ball is fired downwards

OpenStudy (anonymous):

and wrt to the sapceship.. it would be moving with a apeed of 150m/s downward

OpenStudy (shane_b):

Another way to look at it would be to picture the ball and ship moving upward at 100 m/s. The ball is then shot down while it's already moving upwards at 100 m/s. The shot itself is just providing enough force to slow it down by 50m/s. The sum would be 100+(-50)=50...AS @dominusscholae said.

OpenStudy (experimentx):

hint: draw picture ..

OpenStudy (anonymous):

vbm = vb - vm == vb - 0 = 50m/s...

OpenStudy (anonymous):

i made a mistake...becoz...i took vs wrt to vb

ganeshie8 (ganeshie8):

to a static observer, spaceship moves @ 100 m/s ball moves in the same direction as spaceship so, add the velocity : initial velocity 100 + 50 = 150 m/s

OpenStudy (shane_b):

The only way an observer on the ground would see the ball moving at 150 m/s is if it were actually shot UP from the spaceship (which is already going at +100 m/s) at a rate of +50 m/s. The question states the ball is shot down...which is why it gives a rate of -50 m/s.

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):